(3.1) Find an expression for 12 || ~ u + ~ v || 2 + 12 || ~ u − ~ v || 2 in terms of || ~ u || 2 + || ~ v || 2

(3.2) Find an expression for || ~ u + ~ v || 2 − || ~ u − ~ v || 2 in terms of ~ u · ~ v

3.3) Use the result of (3.2) to deduce an expression for || ~ u + ~ v || 2 whenever ~ u and ~ v are orthogonal

to each other.


1
Expert's answer
2021-07-18T17:05:44-0400

(3.1). Given arbitrary vectors "u" and "v." We get


"||u+v||^2=(u+v)(u+v)""=(u,u)+(u,v)+(v,u)+(v,v)""=||u||^2+2(u,v)+||v||^2"




"||u-v||^2=(u-v)(u-v)""=(u,u)-(u,v)-(v,u)+(v,v)""=||u||^2-2(u,v)+||v||^2"

Then



"||u+v||^2+||u-v||^2""=||u||^2+2(u,v)+||v||^2+||u||^2-2(u,v)+||v||^2""=2||u||^2+2||v||^2"

Therefore



"\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u-v||^2=||u||^2+||v||^2"

(3.2).



"||u+v||^2-||u-v||^2""=||u||^2+2(u,v)+||v||^2-||u||^2+2(u,v)-||v||^2""=4(u,v)"

Therefore



"||u+v||^2-||u-v||^2=4u\cdot v"



(3.3). If u and v are orthogonal to each other, then "u\cdot v=0."

Hence


"||u+v||^2-||u-v||^2=0""=>||u+v||^2=|u-v||^2"




"\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u-v||^2""=\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u+v||^2""=||u+v||^2""=\dfrac{1}{2}||u-v||^2+\dfrac{1}{2}||u-v||^2""=||u-v||^2""=||u||^2+||v||^2""||u-v||^2=||u+v||^2=||u||^2+||v||^2"